• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 00:01:58 on October 2, 1998

    Entry number 2093

    keyword=Beam Time Accounting
    Hall A

    10/1/1998 1600-2400
    Experiment #: E94-010
    Shift Leader: R. Michaels
    Shift Workers: P. Zolnierczuk, T. Averett

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    0.5 0.3 1 1 1 1 1 0.7 6.5
    Beam Available
    but Not in Use
    0.5 0.7 0 0 0 0 0 0 1.2
    Beam Not Available
    or Unacceptable
    0 0 0 0 0 0 0 0.3 0.3
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    0.5 0.3 1 1 1 1 1 1 6.8
    Planned Config-
    uration Change
    0 0 0 0 0 0 0 0 0
    Unplanned
    Experiment Down
    0.5 0.7 0 0 0 0 0 01.2
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0.5 0 0 0 0 0 0 0 0.5

    Comments: Shift Summary: Start with Moller tests. Lots of problems - no Moller signal, or very little. 19:20 we stop trying with Moller. They think there may be damage to part of their lead glass detector. 19:30 tests of helicity correlated charge asymmetry -- it is -0.5%, similar to last night. M. Poelker tries to reduce this. Finally inserting half waveplate made it +0.1%. Continue with He3 parallel p=1.484 GeV kinematics until end of shift.

    Data Entry by: R. Michaels